# Limits to extreme event forecasting in chaotic systems

### v1

@@ 4 Conclusions
In this study, we have used the information-theoretic Fano’s inequality to derive the lower bound for predicting extreme events in chaotic time series. The applications discussed here demonstrate the utility of the lower bound obtained (Eq. 5) as a tool for investigating the intrinsic limitations to forecast extreme events in chaotic systems. We have shown that this bound maintains its validity irrespective of the chosen modeling method and plays a crucial role in assessing whether models are functioning near their theoretical optimum. Furthermore, the bound is applicable to scenarios involving both balanced and unbalanced risks and can be used to investigate sources of error from initial conditions, unobserved variables, and suboptimal modeling assumptions (Eq. 6).
The use of Eq. (5) extends beyond the extreme events prediction presented here; it is broadly applicable to any binary classification of events, regardless of whether they are categorized as extreme/non-extreme. Equation (5) also lays the foundations for future extensions to forecasting continuous time signals by employing generalized versions of Fano’s inequality.
@@ 7 Acknowledgments
Appendix A Proof of information-theoretic bound
@@
$P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})=h^{-1}\left(H^{c}(E\mid\boldsymbol{\hat{Q}}^{-})\right)\leq P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f}).$ (12)
@@
where we assume that $21/c^{-}+1/c^{+}=$ 2 to ensure that
$P_{e}^{e,\min}(\boldsymbol{\hat{Q}}^{-})\leq\min\left\{c^{-}P(E=1),c^{+}\left(1-P(E=1)\right)\right\}\leq\frac{c^{-}c^{+}}{c^{+}+c^{-}}=\frac{1}{2}.$ (14)
For each specific state of $\boldsymbol{\hat{Q}}^{-}$, the minimum probability of error is determined by
@@
$P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f})\geq\sum_{\boldsymbol{\hat{q}}^{-}}I(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})=\mathbb{E}[I(\boldsymbol{\hat{Q}}^{-})].$ (17)
Given the concave function $h(p)=-p\log_{2}p-(1-p)\log_{2}(1-p)$ for $p\in[0,0.5]$, the Jensen’s inequality for the random variable $I(\boldsymbol{\hat{Q}}^{-})$ results in
$h\left(\mathbb{E}[I(\boldsymbol{\hat{Q}}^{-})]\right)\geq\mathbb{E}\left[h\left(I(\boldsymbol{\hat{Q}}^{-})\right)\right].$ (18)
On the left hand side of Eq. (18), noticing that $h(p)$ is a non-decreasing function in the range $[0,0.5]$ and utilizing Eq. (17),
$h\left(\mathbb{E}[I(\boldsymbol{\hat{Q}}^{-})]\right)\leq h\left(P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f})\right).$ (19)
@@
$\mathbb{E}\left[h\left(I(\boldsymbol{\hat{Q}}^{-})\right)\right]=\sum_{\boldsymbol{\hat{q}}^{-}}h\left(I(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})$ (20)
$=\sum_{\boldsymbol{\hat{q}}^{-}}h\left(\min\left\{c^{-}P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}),c^{+}(1-P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}))\right\}\right)P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})$
$=\sum_{\boldsymbol{\hat{q}}^{-}}h^{c}\left(P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})=H^{c}(E\mid\boldsymbol{\hat{Q}}^{-}),$
where $H^{c}(E\mid\boldsymbol{\hat{Q}}^{-})$ is defined as the cost-sensitive conditional entropy, and the cost-sensitive binary entropy function is given by
$h^{c}(p)=\begin{cases}h\left(c^{-}p\right)&\text{ for }p\in\left[0,\frac{1}{2c^{-}}\right),\\ h\left(c^{+}(1-p)\right)&\text{ for }p\in\left[\frac{1}{2c^{-}},1\right],\end{cases}$ (21)
@@
$h^{c}(p)=h\left(\min\left\{c^{-}p,c^{+}\left(1-p\right)\right\}\right).$ (22)
@@
$H^{c}(E\mid\boldsymbol{\hat{Q}}^{-})=\mathbb{E}\left[h\left(I(\boldsymbol{\hat{Q}}^{-})\right)\right]\leq h\left(\mathbb{E}[I(\boldsymbol{\hat{Q}}^{-})]\right)\leq h\left(P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f})\right),$ (23)
which implies
@@
$P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})=h^{-1}\left(H^{c}(E\mid\boldsymbol{\hat{Q}}^{-})\right)\leq P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f}).$ (24)

### v2

@@ 4 Conclusions
In this study, we have derived the minimum cost-sensitive probability of error in extreme event forecasting (Eq. 7) along with its information-theoretic lower and upper bounds (Eq. 11). The bounds are rooted in the cost-sensitive Fano’s and Hellman’s inequalities for the Rényi entropy. Furthermore, the minimum probability of error and its bounds are applicable to scenarios involving both balanced and unbalanced risks. The results are also connected to Takens’ embedding theorem using the information can’t hurt inequality, which shows that incorporating additional time lags into the vector of observables can decrease (but never increase) the minimum probability of error. The probability of error for a forecasting model was also decomposed into three sources: uncertainty in the initial conditions, hidden variables, and suboptimal modeling assumptions (Eq. 13).
We have demonstrated the application of these bounds to determine the limits of extreme event prediction in two cases related to fluid dynamics: the Rössler system and the Kolmogorov flow. These applications illustrate the utility of the minimum probability of error and its lower and upper bounds as tools for investigating the intrinsic limitations of forecasting extreme events in chaotic systems. We have shown that the minimum error and its bounds maintain their validity irrespective of the chosen modeling method and play a crucial role in assessing whether models are functioning near their theoretical optimum. Future efforts will be devoted to understanding the limits of predictability for extreme events such as dissipation, wall-shear stress, and wall-pressure in turbulent flows at high Reynolds numbers, which are relevant for advancing the field of external aerodynamics. Nonetheless, the method presented here is generally applicable to problems in other fields such as economics, biology, and finance, among others.
Finally, the use of Eq. (7) and Eq. (11) extends beyond the extreme event predictions presented here; they are broadly applicable to any binary classification of events, whether they are categorized as extreme or non-extreme. Eq. (7) and Eq. (11) also lay the foundation for future extensions to forecasting continuous in time signals by employing generalized versions of Fano’s and Hellman’s inequalities.
@@ 7 Acknowledgments
Appendix A Proof of information-theoretic bounds
@@
$P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})=\mathbb{E}[I(\boldsymbol{\hat{Q}}^{-})],$ (19)
and it is lower and upper bounded as
$h_{\alpha}^{-1}\left(H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})\right)\leq P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})\leq\min\left\{\frac{1}{2}H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-}),C\right\},$ (20)
where $h_{\alpha}(p)$ is a concave function for $p\in[0,1/2]$ when $0<\alpha\leq 2$, and
$C=\min\left\{c^{-}P(E=1),c^{+}(1-P(E=1))\right\}.$
@@
where the factors $c^{-}$ and $c^{+}$ are scaled as $21/c^{-}+1/c^{+}=$ 2 to ensure that
$P^{c}_{e,\min}(\boldsymbol{\hat{Q}}^{-})$$\leq\begin{cases}c^{-}P(E=1),&\text{if }\hat{E}=0\\ c^{+}P(E=0),&\text{if }\hat{E}=1\end{cases}$ (22)
$\leq\min\left\{c^{-}P(E=1),c^{+}\left(1-P(E=1)\right)\right\}$
$=C\leq\frac{c^{-}c^{+}}{c^{+}+c^{-}}=\frac{1}{2}.$
This convention was adopted to avoid $P^{c}_{e,\min}(\boldsymbol{\hat{Q}}^{-})>1/2$, as in those situations, the model with the minimum probability of error could be obtained by flipping the model with $P^{c}_{e,\min}(\boldsymbol{\hat{Q}}^{-})>1/2$ to the one with a probability of $21-P^{c}_{e,\min}(\boldsymbol{\hat{Q}}^{-})\leq 1/$ 2. For each specific state of $\boldsymbol{\hat{Q}}^{-}$, the minimum probability of error is determined by
@@
$I(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})=\min\left\{c^{-}P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}),c^{+}\left(1-P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)\right\}.$ (24)
@@
$P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})=\sum_{\boldsymbol{\hat{q}}^{-}}I(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})=\mathbb{E}[I(\boldsymbol{\hat{Q}}^{-})]\leq P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f}),$ (25)
which is the Bayes error rate typically discussed in statistical classification duda1973 but applied here in the context of extreme event prediction.
Given the concave function
$h_{\alpha}(p)=\lim_{\gamma\rightarrow\alpha}\frac{1}{1-\gamma}\log_{2}\left(p^{\gamma}+(1-p)^{\gamma}\right)$
for $p\in[0,0.5]$ben1978 when $\alpha\in(0,2]$, the Jensen’s inequality for the random variable $I(\boldsymbol{\hat{Q}}^{-})$ results in
$h_{\alpha}\left(P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})\right)=h_{\alpha}\left(\mathbb{E}[I(\boldsymbol{\hat{Q}}^{-})]\right)\geq\mathbb{E}\left[h_{\alpha}\left(I(\boldsymbol{\hat{Q}}^{-})\right)\right].$ (26)
@@
$\mathbb{E}\left[h_{\alpha}\left(I(\boldsymbol{\hat{Q}}^{-})\right)\right]=\sum_{\boldsymbol{\hat{q}}^{-}}h_{\alpha}\left(I(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})$ (27)
$=$ $\sum_{\boldsymbol{\hat{q}}^{-}}h_{\alpha}\left(\min\left\{c^{-}P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}),c^{+}(1-P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}))\right\}\right)$
$P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})$
$=$ $\sum_{\boldsymbol{\hat{q}}^{-}}h_{\alpha}^{c}\left(P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)P(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})=H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-}),$
where $H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})$ is defined as the cost-sensitive conditional entropy, and the cost-sensitive binary entropy function is given by
$h_{\alpha}^{c}(p)=\begin{cases}h_{\alpha}\left(c^{-}p\right)&\text{ for }p\in\left[0,\frac{1}{2c^{-}}\right),\\ h_{\alpha}\left(c^{+}(1-p)\right)&\text{ for }p\in\left[\frac{1}{2c^{-}},1\right],\end{cases}$ (28)
@@
$h_{\alpha}^{c}(p)=h_{\alpha}\left(\min\left\{c^{-}p,c^{+}\left(1-p\right)\right\}\right).$ (29)
The function $h_{\alpha}^{c}(p)$ emerges naturally as a measure of information. The factors $c^{+}$ and $c^{-}$ weight the importance (risk) of each event, whereas the order $\alpha$ controls how the different probabilities in the distribution contribute to the overall measure of uncertainty. For increasing values of $\alpha$, the measure gives more weight to larger probabilities. For example, consider the process of tossing a coin where both heads and tails are assigned equal importance ($c^{+}$ and $c^{-}$ for heads and tails, respectively, with $c^{+}=c^{-}=1$). The greatest amount of information (i.e., uncertainty) regarding the outcome corresponds to the probability $p=1/2$ that maximizes the binary entropy function $h_{\alpha}(p)$. However, when the importance of the outcomes differs (e.g., heads is preferred over tails, $c^{+}>c^{-}$), the greatest uncertainty is achieved at $p=\frac{c^{+}}{c^{+}+c^{-}}$, maximizing cost-sensitive binary entropy $h_{\alpha}^{c}(p)$. In conclusion,
$H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})=\mathbb{E}\left[h_{\alpha}\left(I(\boldsymbol{\hat{Q}}^{-})\right)\right]\leq h_{\alpha}\left(\mathbb{E}[I(\boldsymbol{\hat{Q}}^{-})]\right)=h_{\alpha}\left(P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})\right),$ (30)
which implies
$h_{\alpha}^{-1}\left(H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})\right)\leq P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})=\mathbb{E}\left[I(\boldsymbol{\hat{Q}}^{-})\right].$ (31)
On the other hand, given the binary entropy function $h_{\alpha}(p)$ for $p\in[0,0.5],\alpha\in(0,2]$, it is straight forward to show that
$h_{\alpha}\left(I(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)\geq 2I(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}).$ (32)
Applying the expectation operator to each side of the inequality,
$H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})=\mathbb{E}\left[h_{\alpha}\left(I(\boldsymbol{\hat{Q}}^{-})\right)\right]\geq 2\mathbb{E}\left[I(\boldsymbol{\hat{Q}}^{-})\right]=2P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-}).$ (33)
@@
$h_{\alpha}^{-1}\left(H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})\right)\leq P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})=\mathbb{E}\left[I(\boldsymbol{\hat{Q}}^{-})\right]\leq\min\left\{\frac{1}{2}H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-}),C\right\}$ (34)
Appendix B Proof of inequality of minimum probability of error for additional time lags
A consequence of incorporating additional time lags into the vector observable variables is that
$P_{e,\text{min,LB}}^{c}(\boldsymbol{\hat{Q}}^{-l})\leq P_{e,\text{min,LB}}^{c}(\boldsymbol{\hat{Q}}^{-p}),\ \text{for}\ l>p,$ (35)
$P_{e,\text{min,UB}}^{c}(\boldsymbol{\hat{Q}}^{-l})\leq P_{e,\text{min,UB}}^{c}(\boldsymbol{\hat{Q}}^{-p}),\ \text{for}\ l>p,$
where $l$ and $p$ denote the number of time lags in $\boldsymbol{\hat{Q}}^{-l}$ and $\boldsymbol{\hat{Q}}^{-p}$, respectively, i.e.,
$\boldsymbol{\hat{Q}}^{-l}$ $=[\boldsymbol{\hat{Q}}(t),\boldsymbol{\hat{Q}}(t-\delta t_{1}),\dots,\boldsymbol{\hat{Q}}(t-\delta t_{l})]\pm\delta\boldsymbol{\hat{Q}}^{-l},$ (36)
$\boldsymbol{\hat{Q}}^{-p}$ $=[\boldsymbol{\hat{Q}}(t),\boldsymbol{\hat{Q}}(t-\delta t_{1}),\dots,\boldsymbol{\hat{Q}}(t-\delta t_{p})]\pm\delta\boldsymbol{\hat{Q}}^{-p}.$
Proof.
First, we prove the cost-sensitive conditional entropy inequality,
$H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-l})\leq H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-p}).$ (37)
Noting that $\boldsymbol{\hat{Q}}^{-l}=[\boldsymbol{\hat{Q}}^{-p},\boldsymbol{\hat{Q}}^{-r}]$, where $\boldsymbol{\hat{Q}}^{-r}=[\boldsymbol{\hat{Q}}(t-\delta t_{p+1}),\dots,\boldsymbol{\hat{Q}}(t-\delta t_{l})]\pm\delta\boldsymbol{\hat{Q}}^{-r}$, proving Eq (37) is equivalent to proving
$H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-p},~{}\boldsymbol{\hat{Q}}^{-r})\leq H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-p}).$ (38)
By the law of total probability applied to the conditional probability:
$P(E=1\mid\boldsymbol{\hat{Q}}^{-p}=\boldsymbol{\hat{q}}^{-p})$ (39)
$=$ $\sum_{\boldsymbol{\hat{q}}^{-r}}P(E=1\mid\boldsymbol{\hat{Q}}^{-p}=\boldsymbol{\hat{q}}^{-p},~{}\boldsymbol{\hat{Q}}^{-r}=\boldsymbol{\hat{q}}^{-r})P(\boldsymbol{\hat{Q}}^{-r}=\boldsymbol{\hat{q}}^{-r}\mid\boldsymbol{\hat{Q}}^{-p}=\boldsymbol{\hat{q}}^{-p}).$
The cost-sensitive entropy function $h_{\alpha}^{c}$ is a concave function when $\alpha\in(0,2]$, applying Jensen’s inequality,
$h_{\alpha}^{c}\left(P(E=1\mid\boldsymbol{\hat{Q}}^{-p}=\boldsymbol{\hat{q}}^{-p})\right)$ (40)
[…]
$\geq$ $\sum_{\boldsymbol{\hat{q}}^{-r}}h_{\alpha}^{c}\left(P(E=1\mid\boldsymbol{\hat{Q}}^{-p}=\boldsymbol{\hat{q}}^{-p},~{}\boldsymbol{\hat{Q}}^{-r}=\boldsymbol{\hat{q}}^{-r})\right)P(\boldsymbol{\hat{Q}}^{-r}=\boldsymbol{\hat{q}}^{-r}\mid\boldsymbol{\hat{Q}}^{-p}=\boldsymbol{\hat{q}}^{-p}).$
Applying the inequality to the right hand side of Eq. (37), we get
$H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-p})$ (41)
$=$ $\sum_{\boldsymbol{\hat{q}}^{-p}}h_{\alpha}^{c}\left(P(E=1\mid\boldsymbol{\hat{Q}}^{-p}=\boldsymbol{\hat{q}}^{-p})\right)P(\boldsymbol{\hat{Q}}^{-p}=\boldsymbol{\hat{q}}^{-p})$
[…]
$P(\boldsymbol{\hat{Q}}^{-r}=\boldsymbol{\hat{q}}^{-r}\mid\boldsymbol{\hat{Q}}^{-p}=\boldsymbol{\hat{q}}^{-p})P(\boldsymbol{\hat{Q}}^{-p}=\boldsymbol{\hat{q}}^{-p})$
$=$ $\sum_{\boldsymbol{\hat{q}}^{-p}}\sum_{\boldsymbol{\hat{q}}^{-r}}h_{\alpha}^{c}\left(P(E=1\mid\boldsymbol{\hat{Q}}^{-p}=\boldsymbol{\hat{q}}^{-p},~{}\boldsymbol{\hat{Q}}^{-r}=\boldsymbol{\hat{q}}^{-r})\right)P(\boldsymbol{\hat{Q}}^{-r}=\boldsymbol{\hat{q}}^{-r},~{}\boldsymbol{\hat{Q}}^{-p}=\boldsymbol{\hat{q}}^{-p})$
$=$ $H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-p}~{},\boldsymbol{\hat{Q}}^{-r})=H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-l}).$
Finally, the conditional entropy inequality is applied to the information theoretic-bounds
$P_{e,\text{min,LB}}^{c}(\boldsymbol{\hat{Q}}^{-l})=h_{\alpha}^{-1}\left(H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-l})\right),$ (42)
$P_{e,\text{min,LB}}^{c}(\boldsymbol{\hat{Q}}^{-p})=h_{\alpha}^{-1}\left(H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-p})\right),$
$P_{e,\text{min,UB}}^{c}(\boldsymbol{\hat{Q}}^{-l})=\min\left\{\frac{1}{2}H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-}),C\right\},$
$P_{e,\text{min,UB}}^{c}(\boldsymbol{\hat{Q}}^{-p})=\min\left\{\frac{1}{2}H_{\alpha}^{c}(E\mid\boldsymbol{\hat{Q}}^{-}),C\right\},$
by taking into account that $P^{c}_{e,\min}(\boldsymbol{\hat{Q}}^{-p})\leq 1/2$,
$P_{e,\text{min,LB}}^{c}(\boldsymbol{\hat{Q}}^{-l})\leq P_{e,\text{min,LB}}^{c}(\boldsymbol{\hat{Q}}^{-p}),\ \text{for}\ l>p,$ (43)
$P_{e,\text{min,UB}}^{c}(\boldsymbol{\hat{Q}}^{-l})\leq P_{e,\text{min,UB}}^{c}(\boldsymbol{\hat{Q}}^{-p}),\ \text{for}\ l>p,$
Appendix C Continuous extension of information-theoretic bound
We show that the bound for minimum cost-sensitive probability of error also holds for continuous observables just by replacing the cost-sensitive entropy $H_{\alpha}^{c}$ by the cost-sensitive entropy $\mathcal{H}_{\alpha}^{C}$ conditioned on a continuous variable. The minimum cost-sensitive probability of error achievable by any model with continuous $\boldsymbol{\hat{Q}}^{-}$ is
@@ Appendix C Continuous extension of information-theoretic bound
$h_{\alpha}^{-1}\left(\mathcal{H_{\alpha}}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})\right)\leq P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})=\mathbb{E}[\mathcal{I}(\boldsymbol{\hat{Q}}^{-})]\leq\min\left\{\frac{1}{2}\left(\mathcal{H_{\alpha}}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})\right),C\right\}.$ (44)
The cost-sensitive conditional entropy is defined as
$\mathcal{H_{\alpha}}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})$ $=\int_{\boldsymbol{\hat{Q}}^{-}}h_{\alpha}^{c}\left(P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)\rho_{\boldsymbol{\hat{Q}}^{-}}(\boldsymbol{\hat{q}}^{-})\mathrm{d}\boldsymbol{\hat{q}}^{-},$ (45)
where $\rho_{\boldsymbol{\hat{Q}}^{-}}(\boldsymbol{\hat{q}}^{-})$ is the probability density function of $\boldsymbol{\hat{Q}}^{-}$ and the integral is conducted over the support of $\boldsymbol{\hat{Q}}^{-}$. The minimum probability of error at each value is defined as
$\mathcal{I}(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})=\min\left\{c^{-}P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-}),c^{+}\left(1-P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)\right\}.$ (46)
Proof.
Similar to Eq. (21), the overall cost-sensitive probability of error is a weighted integral of the probability of error for each value of $\boldsymbol{\hat{Q}}^{-}$,
$P_{e}^{c}(\boldsymbol{\hat{Q}}^{-},\hat{f})$ $=\int_{\boldsymbol{\hat{Q}}^{-}}\biggl{(}c^{-}P(\hat{E}=0,E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})$ (47)
$\qquad+c^{+}P(\hat{E}=1,E=0\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\biggr{)}\rho_{\boldsymbol{\hat{Q}}^{-}}(\boldsymbol{\hat{q}}^{-})\mathrm{d}\boldsymbol{\hat{q}}^{-},$
Applying Eq. (47) and Eq. (23) we have,
$P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})=\int_{\boldsymbol{\hat{Q}}^{-}}\mathcal{I}(\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\rho_{\boldsymbol{\hat{Q}}^{-}}(\boldsymbol{\hat{q}}^{-})\mathrm{d}{\boldsymbol{\hat{q}}^{-}}=\mathbb{E}[\mathcal{I}(\boldsymbol{\hat{Q}}^{-})].$ (48)
Analogous to Eq. (30),
$\mathbb{E}\left[h_{\alpha}\left(\mathcal{I}(\boldsymbol{\hat{Q}}^{-})\right)\right]\leq h_{\alpha}\left(\mathbb{E}[\mathcal{I}(\boldsymbol{\hat{Q}}^{-})]\right)=h_{\alpha}\left(P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})\right),$ (49)
where
$\mathbb{E}\left[h_{\alpha}\left(I(\boldsymbol{\hat{Q}}^{-})\right)\right]$ $=\int_{\boldsymbol{\hat{Q}}^{-}}h_{\alpha}\left(I(\boldsymbol{\hat{q}}^{-})\right)\rho_{\boldsymbol{\hat{Q}}^{-}}(\boldsymbol{\hat{q}}^{-})\mathrm{d}\boldsymbol{\hat{q}}^{-}$ (50)
$=\int_{\boldsymbol{\hat{Q}}^{-}}h_{\alpha}^{c}\left(P(E=1\mid\boldsymbol{\hat{Q}}^{-}=\boldsymbol{\hat{q}}^{-})\right)\rho_{\boldsymbol{\hat{Q}}^{-}}(\boldsymbol{\hat{q}}^{-})\mathrm{d}\boldsymbol{\hat{q}}^{-}$
$=\mathcal{H_{\alpha}}^{c}(E\mid\boldsymbol{\hat{Q}}^{-}).$
On the other hand, similar to Eq. (33),
$\mathcal{H_{\alpha}}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})=\mathbb{E}\left[h_{\alpha}\left(\mathcal{I}(\boldsymbol{\hat{Q}}^{-})\right)\right]\geq 2\mathbb{E}\left[\mathcal{I}(\boldsymbol{\hat{Q}}^{-})\right]=2P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-}).$ (51)
In conclusion,
@@
$h_{\alpha}^{-1}\left(\mathcal{H_{\alpha}}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})\right)\leq P_{e,\text{min}}^{c}(\boldsymbol{\hat{Q}}^{-})=\mathbb{E}\left[\mathcal{I}(\boldsymbol{\hat{Q}}^{-})\right]\leq\min\left\{\frac{1}{2}\left(\mathcal{H_{\alpha}}^{c}(E\mid\boldsymbol{\hat{Q}}^{-})\right),C\right\}.$ (52)
Appendix D Results for higher thresholds and confusion matrix for $\hat{f}^{DT}$
To address the sensitivity of the results to the intensity of the extreme events in the Rössler system and the Kolmogorov flow, we repeated the analysis with higher threshold values in both cases. The minimum probability of error is shown in Fig. 7 and Fig. 8 as a function of $\delta t$ for the cases discussed above. The main observation is that increasing the threshold for extreme events makes the prediction more challenging, due to the higher scarcity of events, which renders them more unpredictable. Nonetheless, the trends discussed in the main text regarding the behavior of $\delta t$ remain unchanged.
Figure 7: Normalized probability of error for extreme events prediction in the Rössler system for the threshold $\eta=\bar{\theta}_{3}+6\sigma_{\theta_{3}}$ using observable (a) $\hat{Q}^{-}_{1}$ and (b) $\boldsymbol{\hat{Q}}^{-}_{2}$. $\bar{P}_{e,\min}^{c}(\hat{Q}^{-}_{1})$, $\bar{P}_{e,\min}^{c}(\boldsymbol{\hat{Q}}^{-}_{2})$, and $\bar{P}_{e,\min}^{c}(\boldsymbol{Q}^{-})$ are the minimum probability of error using the observable $\hat{Q}^{-}_{1}$, $\boldsymbol{\hat{Q}}^{-}_{2}$, and $\boldsymbol{Q}^{-}$, respectively.
Figure 8: Normalized, cost-sensitive minimum probability of error for extreme event prediction in the Kolmogorov flow for $\eta=\bar{D}+3\sigma_{D}$.
The confusion matrix for the decision tree model to predict extreme events in the Rössler system using observable $\hat{Q}^{-}_{1}$ is shown in Table 1 for $\delta t=0.075$.
Decision Tree $\hat{E}=0$ $\hat{E}=1$
$E=0$ 0.9716 0.0060
$E=1$ 0.0030 0.0194
Table 1: Confusion Matrix of Decision Tree model. The normalized probability of error of the model can be calculated as $\bar{P}_{e}^{c}(\hat{Q}^{-}_{1},\hat{f}^{DT};\delta t=0.075)=\frac{0.0060+0.0030}{0.0030+0.0194}=0.3554$.